Circular No. 7187

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cf IN UGC 8539 S. Benetti, Telescopio Nazionale Galileo, communicates: “Inspection

of a preliminarily reduced CCD spectrogram (range 330–756 nm, resolution 0.25 nm) obtained on June 4.04 UT with the William Herschel Telescope (+ ISIS) at La Palma confirms that SN 1999cf (cf. IAUC 7178) is a type-Ia supernova at maximum light. The spectrum is dominated by P-Cyg lines of intermediate-mass elements superimposed on a blue continuum. Visible lines include Ca II (H and K), Si II (597.2 and 635.5 nm), and S II (546.8, 561.2, 565.4 nm). The expansion velocities deduced from the minima of Si II 635.5-nm and Ca II H and K lines, corrected by the redshift (7353 km/s, from NED) of the parent galaxy, are ∼ 12 000 and 15 800 km/s, respectively.”

XTE J1550–564 R. Jain, C. Bailyn, and J. Greene, Yale University; J. Orosz, Penn-

sylvania State University; J. McClintock, Harvard-Smithsonian Center for Astrophysics; and R. Remillard, Massachusetts Institute of Technology, report YALO consortium observations of XTE J1550–564: “We have obtained data from May 14.40 to June 1.3 UT using the YALO 1-m telescope at Cerro Tololo. The uncalibrated R-band data, with relative precision of 0.02 mag, show a 1-mag decrease during this period; hence the optical light curve has not yet reached the true quiescent level. Between May 14.4 and 23.28, we obtained on average three 1200-s Johnson R exposures per night; each was separated by several hours. We detrended the lightcurve by removing a linear decay of 0.05 mag/day and find clear evidence for a modulation of 0.1 mag (full amplitude). Using the CLEAN algorithm and performing a fourier transform, we find a period of 1.5 ± 0.1 days. Similarly, by using the phase-dispersion-minimization technique, we find a clear period at 1.5 days. If this period of ∼ 1.5 days (which was suggested earlier by Soria and Wu, IAUC 7184) is correct, then the modulation we see may be due to a ‘superhump’ or to an illumination effect. On the other hand, if the light from the secondary dominates the light from the accretion disk (as suggested by the strong absorption lines reported by Soria and Wu), then the modulation we observe could be an ellipsoidal modulation. In this case, the orbital period would be near 3 days. Further spectroscopic observations will be needed to definitively determine the orbital period.”

1999 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green